# Efficient and Reproducible Biomedical Question Answering using Retrieval Augmented Generation

### v1

@@ Abstract
Index Terms:
Biomedical Information Retrieval, Retrieval-Augmented Generation, Hybrid Retrieval, Large Language Models, PubMed, Information Retrieval Systems.
@@ I Introduction
Despite the potential of biomedical RAG systems, existing solutions often suffer from limited scalability, poor reproducibility, and suboptimal retrieval performance on large datasets such as PubMed [^1] https://pubmed.ncbi.nlm.nih.gov/. Existing benchmarks for medical question answering, such as MedExpQA [4] and MIRAGE [5], lack reproducibility and scalable retrieval solutions. Most retrieval methods rely on either sparse bag-of-words vectors such as BM25 [6] or dense vectors created through transformer-based models such as BioBERT [7] and MedCPT [8]. However, hybrid approaches that integrate both techniques remain under-investigated, especially from a system perspective: the inherent trade-offs between retrieval strategies, their indexing and response times, and the resulting generator accuracy are crucial for practical RAG applications and have been largely unexplored. Hybrid retrieval methods combine the strengths of sparse and dense retrieval: a probabilistic retriever (e.g., BM25) efficiently reduces the search space by filtering a large corpus, while a neural reranker (e.g., MedCPT’s cross-encoder) refines document rankings based on semantic relevance. This two-step approach balances computational efficiency and retrieval precision, mitigating the limitations of stand-alone methods. Although hybrid retrieval has been explored in general NLP tasks [9], its application in large-scale biomedical QA remains limited, particularly in real-world implementations. Developing an effective biomedical QA system requires addressing several challenges: (i) Efficient Retrieval at Scale: Processing millions of biomedical documents under reasonable indexing times while maintaining low-latency retrieval; (ii) Relevance Optimization: Improving document ranking by integrating lexical retrieval with neural reranking; (iii) Context Integration: Structuring retrieved documents effectively to generate factually accurate and verifiable responses. This work presents a scalable and reproducible RAG system for biomedical QA, systematically evaluating hybrid retrieval strategies. The key contributions include:
@@ IV Evaluation of the Final System
After selecting the most efficient and effective components for our RAG system, we evaluate the final Retrieval-Augmented Generation (RAG) system with the hybrid retrieval approach on the full 24M document PubMed corpus. for retrieval effectiveness, response time, and answer correctness. We use a Linux server with an Intel Broadwell processor (16 cores, 30GB RAM) and an NVIDIA A30 GPU (24GB VRAM) for cross-encoder reranking.

### v2

@@ Abstract
footnotetext: This paper was published at the IEEE Swiss Conference on Data Science (SDS 2025). DOI: 10.1109/SDS66131.2025.00029.
@@ I Introduction
Despite the potential of biomedical RAG systems, existing solutions often suffer from limited scalability, poor reproducibility, and suboptimal retrieval performance on large datasets such as PubMed. Existing benchmarks for medical question answering, such as MedExpQA [1] and MIRAGE [16], lack reproducibility and scalable retrieval solutions. Most retrieval methods rely on either sparse bag-of-words vectors such as BM25 [12] or dense vectors created through transformer-based models such as BioBERT [7] and MedCPT [5]. However, hybrid approaches that integrate both techniques remain under-investigated, especially from a system perspective: the inherent trade-offs between retrieval strategies, their indexing and response times, and the resulting generator accuracy are crucial for practical RAG applications and have been largely unexplored. Hybrid retrieval methods combine the strengths of sparse and dense retrieval: a probabilistic retriever (e.g., BM25) efficiently reduces the search space by filtering a large corpus, while a neural reranker (e.g., MedCPT’s cross-encoder) refines document rankings based on semantic relevance. This two-step approach balances computational efficiency and retrieval precision, mitigating the limitations of stand-alone methods. Although hybrid retrieval has been explored in general NLP tasks [9], its application in large-scale biomedical QA remains limited, particularly in real-world implementations. Developing an effective biomedical QA system requires addressing several challenges: (i) Efficient Retrieval at Scale: Processing millions of biomedical documents under reasonable indexing times while maintaining low-latency retrieval; (ii) Relevance Optimization: Improving document ranking by integrating lexical retrieval with neural reranking; (iii) Context Integration: Structuring retrieved documents effectively to generate factually accurate and verifiable responses. This work presents a scalable and reproducible RAG system for biomedical QA, systematically evaluating hybrid retrieval strategies. The key contributions include:
@@ IV Evaluation of the Final System
After selecting the most efficient and effective components, we evaluate the final hybrid RAG system on the full 24M-document PubMed corpus for retrieval effectiveness, response time, and answer correctness.